callmeizzy1234:

What is the result of subtracting (3x^2 - 2x) from its conjugate?

callmeizzy1234:

note this is not a problem for school

jhonyy9:

like a first step - what is a conjugate ?

jhonyy9:

@supie

callmeizzy1234:

but i have a test coming up to skip the grade

callmeizzy1234:

to combine like trems

callmeizzy1234:

@jhonyy9

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 like a first step - what is a conjugate ? \(\color{#0cbb34}{\text{End of Quote}}\) you need to know what is a conjugate ...

callmeizzy1234:

well i gtg can we resume tomorrow

callmeizzy1234:

i need to know all this by the end of the week to get high school credits

jhonyy9:

so just you need to know what mean ,,the conjugate of ... " ? @XioGonz do you know ?

jhonyy9:

have you learn about conjugate ?

jhonyy9:

no ?

XioGonz:

Conjugate--change the sign(+ to −, or − to +) in the middle of two terms.

jhonyy9:

exactly - so how will be in case of this posted problem ?

jhonyy9:

what is the conjugate of 3x^2 -2x ?

jhonyy9:

use pls what you ve wrote above

XioGonz:

I am able to define it but I am unaware of how to solve this problem either - I apologize

jhonyy9:

why ? you ve said Conjugate--change the sign(+ to −, or − to +) in the middle of two terms. so than pls change the sign of - in + and subtract the 3x^2 -2x from its conjugate this is so easy

jhonyy9:

@XioGonz pls

jhonyy9:

courage - you are on a right way

XioGonz:

I've helped with what I can but I'm assuming it goes somewhat like this : (3x^2 + 2x)-(3x^2 - 2x) =4x ?

XioGonz:

@jhonyy9

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz I've helped with what I can but I'm assuming it goes somewhat like this : (3x^2 + 2x)-(3x^2 - 2x) =4x ? \(\color{#0cbb34}{\text{End of Quote}}\) correct - good job - congrats !

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz I've helped with what I can but I'm assuming it goes somewhat like this : (3x^2 + 2x)-(3x^2 - 2x) =4x ? \(\color{#0cbb34}{\text{End of Quote}}\) correct - good job - congrats ! \(\color{#0cbb34}{\text{End of Quote}}\) Thank you!

jhonyy9:

yw